40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 11 04:08:53 2015, Jenne, Update, LSC, New Locking Paradigm? 
    Reply  Wed Feb 11 22:13:44 2015, Jenne, Update, LSC, New Locking Paradigm - LSC model changes 
       Reply  Thu Feb 12 01:43:09 2015, rana, Update, LSC, New Locking Paradigm - LSC model changes TF.gif
       Reply  Thu Feb 12 11:14:29 2015, Jenne, Update, LSC, New Locking Paradigm - Loop-gebra ALS_REFL_comboLockingCartoon_11Feb2015.PDFALS_REFL_comboLocking_11Feb2015.zip
          Reply  Thu Feb 12 11:59:58 2015, Koji, Update, LSC, New Locking Paradigm - Loop-gebra 
             Reply  Thu Feb 12 19:18:49 2015, Jenne, Update, LSC, New Locking Paradigm - Loop-gebra 
                Reply  Fri Feb 13 03:28:34 2015, rana, Update, LSC, New Locking Paradigm - Loop-gebra 
                   Reply  Sat Feb 14 20:20:24 2015, Jenne, Update, LSC, ALS fool cartoon ALSfool_LoopDiagram.pngALSfool_LoopDiagram.graffle.zip
                      Reply  Tue Feb 17 04:04:32 2015, Jenne, Update, LSC, ALS fool math OpenLoopGainComparison_16Feb2015.pngCancellationTFs_DecouplingIsArefl.png
                         Reply  Tue Feb 17 16:36:08 2015, Jenne, Update, LSC, ALS fool math ALSfool_LoopDiagram_17Feb2015.pngALS_REFL_comboLocking_16Feb2015.nb.zip
       Reply  Thu Feb 12 22:28:16 2015, Jenne, Update, LSC, New Locking Paradigm - LSC model changes, screens modified 
          Reply  Sat Mar 7 19:15:17 2015, Jenne, Update, LSC, Modified zero crossing triggering LatchLogic_6March2015.png
    Reply  Thu Feb 12 03:43:54 2015, ericq, Update, LSC, 3F PRMI at zero ALS CARM Feb12_negativeTR.png
Message ID: 11016     Entry time: Thu Feb 12 19:18:49 2015     In reply to: 11012     Reply to this: 11020
Author: Jenne 
Type: Update 
Category: LSC 
Subject: New Locking Paradigm - Loop-gebra 

EDIT, JCD, 17Feb2015:  Updated loop diagram and calculation: http://131.215.115.52:8080/40m/11043


Okay, Koji and I talked (after he talked to Rana), and I re-looked at the original cartoon from when Rana and I were thinking about this the other day.

The original idea was to be able to actuate on the MC frequency (using REFL as the sensor), without affecting the ALS loop.  Since actuating on the MC will move the PSL frequency around, we need to tell the ALS error signal how much the PSL moved in order to subtract away this effect. (In reality, it doesn't matter if we're actuating on the MC or the ETMs, but it's easier for me to think about this way around).  This means that we want to be able to actuate from point 10 in the diagram, and not feel anything at point 4 in the diagram (diagram from http://131.215.115.52:8080/40m/11011)

This is the same as saying that we wanted the green trace in http://131.215.115.52:8080/40m/11009 to be zero.

So.  What is the total TF from 10 to 4? 

{\rm TF}_{\rm (10 \ to \ 4)} = \frac{D_{\rm cpl} + {\color{DarkRed} A_{\rm refl}} {\color{DarkGreen} P_{\rm als}}}{1-{\color{DarkRed} A_{\rm refl} G_{\rm refl} S_{\rm refl} P_{\rm refl}} - {\color{DarkGreen} A_{\rm als} G_{\rm als} S_{\rm als}} ({\color{DarkGreen} P_{\rm als}} + D_{\rm cpl} {\color{DarkRed} G_{\rm refl} P_{\rm refl} S_{\rm refl}})}

So, to set this equal to zero (ALS is immune to any REFL loop actuation), we need D_{\rm cpl} = - {\color{DarkRed} A_{\rm refl}} {\color{DarkGreen} P_{\rm als}}.

Next up, we want to see what this means for the closed loop gain of the whole system.  For simplicity, let's let H_* = A_* G_* S_* P_*, where * can be either REFL or ALS. 

Recall that the closed loop gain of the system (from point 1 to point 2)  is

{\rm TF}_{\rm (1 \ to \ 2)} = \frac{1}{1-{\color{DarkRed} A_{\rm refl} G_{\rm refl} S_{\rm refl} P_{\rm refl}} - {\color{DarkGreen} A_{\rm als} G_{\rm als} S_{\rm als}} ({\color{DarkGreen} P_{\rm als}} + D_{\rm cpl} {\color{DarkRed} G_{\rm refl} P_{\rm refl} S_{\rm refl}})} , so if we let  D_{\rm cpl} = - {\color{DarkRed} A_{\rm refl}} {\color{DarkGreen} P_{\rm als}} and simplify, we get

{\rm TF}_{\rm (1 \ to \ 2)} = \frac{1}{1-{\color{DarkRed} H_{\rm refl}} - {\color{DarkGreen} H_{\rm als}} + {\color{DarkRed} H_{\rm refl}}{\color{DarkGreen} H_{\rm als}}}

This seems a little scary, in that maybe we have to be careful about keeping the system stable.  Hmmmm.  Note to self:  more brain energy here.


Also, this means that I cannot explain why the filter wasn't working last night, with the guess of a complex pole pair at 1Hz for the MC actuator.  The  ALS plant has a cavity pole at ~80kHz, so for our purposes is totally flat.  The only other thing that comes to mind is the delays that exist because the ALS signals have to hop from computer to computer.  But, as Rana points out, this isn't really all that much phase delay below 100Hz where we want the cancellation to be awesome. 

I propose that we just measure and vectfit the transfer function that we need, since that seems less time consuming than iteratively tweaking and checking. 

Also, I just now looked at the wiki, and the MC2 suspension resonance for pos is at 0.97Hz, although I don't suspect that that will have changed anything significantly above a few Hz.  Maybe it makes the cancellation right near 1Hz a little worse, but not well above the resonance.

 

ELOG V3.1.3-